OpenStudy (danysanchy97):

The least common multiple of 3, 4, 6, and 8 is A. 24. B. 8. C. 72. D. 96.

OpenStudy (photon336):

Basically it's the smallest multiple of that number set

OpenStudy (photon336):

it cannot be 8. it has to be greater than all the numbers but the smallest multiple that all the numbers can go into. what do you think?

OpenStudy (danysanchy97):

a? @Photon336

OpenStudy (shansha1):

its A.24

OpenStudy (janu16):

Because 3 is prime, 4=2*2, 6=2*3 and 8=2*2*2 you have to have each prime at least as many times as it is used in any number. So you would need three 2's but only one 3 which gives you 24.

OpenStudy (janu16):

stp giving direct answer plz. explan to them

OpenStudy (janu16):

so they understand